OpenStudy (anonymous):

hey guys! new here, please help Let OP = 2i + 3j and OQ = -6i + 4j. Let R be a point on PQ such that PR : RQ = k : 1, where k > 0. (a) Express OR in terms of k, i and j.

OpenStudy (anonymous):

 hehehe... rely on others take so much time. I try once.

OpenStudy (anonymous):

Let OQ =\(\vec q\) \(OP=\vec p\) \(OR =\vec r\)

OpenStudy (anonymous):

the proportion is \(\dfrac{PR}{RQ}=\dfrac{k}{1}\) that is PR = k RQ

OpenStudy (anonymous):

and we have \(vec {QP}=\vec p-\vec q\)

OpenStudy (anonymous):

\(\vec QR =(1/k)\vec PR\)

OpenStudy (anonymous):

hence we can go from O to R by 

OpenStudy (anonymous):

that is \(\vec q +\vec {QR}\)

OpenStudy (anonymous):

as above, \(\vec{QR}=\dfrac{\vec{QP}}{k}=\dfrac{\vec p-\vec q}{k}\)

OpenStudy (anonymous):

Replace all, we have \(\vec {OR}= \vec q+\dfrac {\vec p-\vec q}{k}\)

OpenStudy (anonymous):

All you need is replace \(\vec q= 2i+3j\) and \(\vec p= -6i+4j\)

OpenStudy (anonymous):

ok, let me do it to save time. I need review also

OpenStudy (anonymous):

\(\vec {OR}=(2i+3j) + \dfrac{(-6i+4j )-(2i+3j)}{k}=(2i+3j) + \dfrac{(-8i+j )}{k}\)

OpenStudy (anonymous):

That is the answer

OpenStudy (anonymous):

um i dont think so because the answer here doesnt match imsp206.netvigator.com/~norme/Main/Maths/AMA/Rev-old.pdf

OpenStudy (anonymous):

[Solution] (a) OR = +1 + k OP kOQ = 2 - 6k k + 1 i + 3 + 4k k + 1 j

OpenStudy (anonymous):

hehehe.. if it is so, I am sorry for my dumb solution.